Write ohm's law.
1
Expert's answer
2020-10-25T18:31:44-0400

For homogeneous part of the circuit

"\displaystyle I = \frac{V}{R}"

where

"V" - voltage

"I" - current

"R" - total resistance of the one or more elements


If the is a source with internal resistance, Ohm's law can be written as

"\displaystyle I = \frac{\varepsilon}{r + R}"

where

"I" - current in the circuit

"r" - internal resistance of the current source

"R" - resistance of the circuit

"\varepsilon" - electromotive force.